OpenStudy (anonymous):

It is my last assignment for today!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! ASAP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! PLEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEASSSSEEEEEE!!!!!!!! HELP!!!!!!!!! 1. Which is the equation of the given line in point-slope form? (The points are, 1 1/2, 2 & -4,-1) A. y+1=6/7 (x+4) B. y+1=6/11 (x+4) C. y+4=11/6 (x+1) D. y-1=6/11(x-4) 2. Which is an equation of the given line in standard form? (The line is, 1 1/2, 2 & -4,-1) A. -11x-6y=-13 B. -6x+11y=13 C. -6x+11y=-13 D. -6x+7y=-17 3. Which is an equation of the given line in standard form? (The line is, -4,-3 & 2,-3) A. x = –3 B. x = 3 C. y = –3 D. y = –3x 4. Which is the equation of the given line in slope-intercept form? (The line is, -1,2 & 1,-4) A. y=-3x-1 B. y=3x+1 C. y=3x-1 D. y=-1/3x-1 5. Which is the equation of the given line in standard form? (The line is, 3,2 & 3,-4) A. x = 3 B. x = –3 C. y = 3 D. y = 3x PLEASE EXPLAIN HOW YOU GOT IT.... Thank you!!! :)

OpenStudy (the_fizicx99):

*Cracks fingers* Kk When dealing with (x,y) coordinates close them after its just (x,y) not (x,y and x,y) it'll be considered wrong. So, to get them into slope-intercept form use the formula: M = Y2-y1/X2-x1 Then once you have your slope, you'll use y-y1=m(x-x1)

OpenStudy (anonymous):

Okay thanks that helps me understand it better... will you give me the answers please. :)

OpenStudy (anonymous):

Hello??

OpenStudy (the_fizicx99):

I'm not allowed to do that, its against the CoC.